A rock takes 2.40 s to hit the ground when it is thrown straight up from the cliff with an initial velocity of 8.17 m/s. How long (in s) would it take to reach the ground if it is thrown straight down with the same speed?
Answer #1

The time duration

for of path of time light - Ton = 2 Vo Tg TOA = 2x 8.17 9.8. TOA = 1.667 sec I so *= p. 40-To A = 2.40-11667 t = 6.73 then ti